OpenStudy (anonymous):

solve: x-6 over 2 = 3

OpenStudy (anonymous):

Does it look like this? \[\frac{ x-6 }{ 2 }=3\]

OpenStudy (anonymous):

Yess.

OpenStudy (anonymous):

Ok, so you want to get x all by itself on one side of the equation. Since we have a fraction over two, the first step we could take is to multiply both sides by 2. What do you get once you do this? It should cancel out the 2 on the left side.

OpenStudy (anonymous):

x-6=6?

OpenStudy (anonymous):

Yep, exactly. So now you just have one step left, which is to add 6 to both sides

OpenStudy (anonymous):

ahhh okay so x=12?

OpenStudy (anonymous):

Yes, exactly, you can check it by plugging 12 in for x in the original equation and making sure you get a true statement

OpenStudy (anonymous):

yess. i got it down to 12-6=6 which is correct.